Q Calculate the moment about the point (1.1.1) of a force of 5 kg acting along the a line AB vector where A and B are the point (2.3.4) and (3.5.6) respectively the distance being measured in meters.

Grade:11

1 Answers

Sandeep Pathak
9 years ago
Let’s assume the point (1,1,1) is O. The moment of Force about O is give by
\vec{\tau} = \vec{OA}\times \vec{F}
Force, F in vector form is given by
\vec{F}=5\times 10\times \hat{AB}=\frac{10}{3}\left(\hat{i}+2\hat{j}+2\hat{k} \right )
\vec{OA}=\left(\hat{i}+2\hat{j}+3\hat{k} \right )
This gives
\vec{\tau}=\frac{10}{3}\left(-2\hat{i}+\hat{j}\right )